OpenStudy (anonymous):

Help Plz will medal Suppose that the functions r and s are defined for all real numbers x as follows. r(x)=x-4 s(x)=4x^2 Write the expressions for (r-s)(x) (r*s)(x) and evaluate (r+s)(3)

OpenStudy (amyhashimoto77):

I'm not too sure on this, but don't you just simply plug it in? Do you have the answer(in the back of the book)? (r-s)(x): =(x-4)-(4x^2) =x-4-4x^2

OpenStudy (amyhashimoto77):

(r*s)(x) =(x-4)(4x^2) =4x^3-16x^2

OpenStudy (anonymous):

i cant remember and no book online

OpenStudy (amyhashimoto77):

(r+s)(3) =3((x-4)+(4x^2)) =3(x-4+4x^2) =3x-12+12x^2

OpenStudy (anonymous):

let you know

OpenStudy (anonymous):

all right except last one

OpenStudy (amyhashimoto77):

ok